Given a regression sum of squares equal to 2373.59 and a total sum of squares equal to 2489.52, the R-square value is? A) .953 B) 1.04 C) .047 D) .873
Answer #1

Solution :

Given that,

Regression sum of squares = 2373.59 and

A total sum of squares = 2489.52

R-square = 1 - (Regression sum of squares / A total sum of squares )

= 1 - (2373.59 / 2489.52)

= 1 - 0.953

= 0.047

R-square value is 0.047 .

Option C)